OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A:y-1=6(x+5) B:y+5=6(x-1) C:y+1=6(x-5) D:y-5=6(x+1)

OpenStudy (midhun.madhu1987):

Point slope of a line is given by: y - y1 = m(x - x1) where m= slope

OpenStudy (midhun.madhu1987):

Plug in the values and find the equation